A wire loop of resistance 10 ohm and radius 10cm is kept in the plane of this paper in a uniform magnetic field B the direction of B is perpendicular to the plane of the page and points out of it and it's magnitude is increasing at the rate of 0.50Ts. Determine the magnitude and direction of the induced current in the loop
1
Expert's answer
2020-03-18T10:25:27-0400


We write the law of electromagnetic induction

"\epsilon_i=\frac{\Delta \Phi} { \Delta t}=\frac{\Delta B \cdot S} { \Delta t}=\frac{\Delta B } { \Delta t}\cdot S"

Where

"\frac{\Delta B } { \Delta t}" -the rate of change of the magnetic field

"S=\pi \cdot r^2" -ring area

Then we write down Ohm’s law

"I_i=\frac{\epsilon_i}{R}=\frac{\Delta B } { \Delta t}\cdot \frac{S}{R}=0.5\cdot \frac{3.14 \cdot 0.1^2}{10}=1.57mA"